OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 8) and a radius of 9.

OpenStudy (anonymous):

Just use a circle general formula:\[(x-a)^2+(y-b)^2=r^2\]where (a,b) is the centre coordinate and r is the radius.

OpenStudy (anonymous):

0-8^2?

OpenStudy (anonymous):

just substitute a=0, b=8 and r=9. x and y stay there as separate terms.

OpenStudy (anonymous):

(x-0)^2+(y-b)^2=9^2

OpenStudy (anonymous):

Almost there: you can make (x-0)^2 = x^2 and don't forget to let b=8.

OpenStudy (anonymous):

cool..thanks!